1) The costs associated with a particular process are expected to be $6,000 per year for five years, beginning three years fr solve three. show your work
Answer #1

Solution 1:

Present worth = $6,000 * Cumulative PV factor at 10% for 3 to 7 periods

= $6,000 * 3.13288 = $18,796

Hence option b is correct.

Solution 2:

Future value = $6,000 * Cumulative FV factor at 10% for 5 periods

= $6,000 * 6.1051 = $36,631.

Hence option A is correct

Solution 3:

Equivalent uniform annual worth = $18,796 /. Cumulative PV factor at 10% for 8 periods

= $18,796 / 5.33493 = $3,522

Hence option c is correct.